OpenStudy (anonymous):

Prove this trig identity. Show all working please. :) Cosx / Secx - Tanx = 1 + Sinx

OpenStudy (anonymous):

\[\frac{ cosx }{ secx-tanx } = 1+ sinx\]

hartnn (hartnn):

write sec x as 1/cos x and tan x as sin x/cos x

hartnn (hartnn):

then write cos^2 x as 1- sin^2x = (1-sin x)(1+sin x)

hartnn (hartnn):

so could u prove it ?

OpenStudy (anonymous):

Nope, I need to refer to my notes but I left my book at home (at school atm)

hartnn (hartnn):

\(\large \frac{cos x}{1/cos x-sin x/cos x}\quad =\frac{cos^2x}{1-sin x}=\frac{(1+sin x)(1-sin x)}{1-sin x}\) now its clear ?

OpenStudy (anonymous):

Yep, the 1-sines cancel. :)

hartnn (hartnn):

\(\huge\frac{cos x}{\frac{1-sin x}{cos x}}=cos x .\frac{cos x}{1-sin x}\)

OpenStudy (anonymous):

Oh right, you're multiplying the reciprocal!

hartnn (hartnn):

yup :)

OpenStudy (anonymous):

One more stupid question, if that's the case what happened to the original 1/cosine ?

OpenStudy (anonymous):

\[\frac{ 1 }{ cosx } and \frac{ sinx }{ cosx }\]

OpenStudy (anonymous):

cosx = cosx/1 so you're left with 1-sine/cosine on the bottom

OpenStudy (anonymous):

\[cosx . \frac{ \cos }{ 1 } - \frac{ cosx }{ sinx } \]

OpenStudy (anonymous):

or is it \[cosx.\frac{ cosx }{ 1 } - \frac{ sinx }{ cosx }\]

OpenStudy (anonymous):

I understand where the cos^2x comes from but there are 3 cosines and I don't know where one of them disappeared to :]

hartnn (hartnn):

sec x = 1/ cos x tan x = sin x / cos x so your denominator is: 1/cos x -sin x/cos x = (1-sin x)/cos x

OpenStudy (anonymous):

Ahh, I see so the common denominator rule is applied, therefore the 2 cosines "merge" ?

hartnn (hartnn):

yup, a/b-c/b=(a-c)/b

OpenStudy (anonymous):

eg \[\frac{ 1 }{ 3 } + \frac{ 2 }{ 3 } = \frac{ 3 }{ 3 }\]

OpenStudy (anonymous):

Awesome! Thank you so much for your patience and time, you've been utmost helpful. :)

OpenStudy (anonymous):

Would it be alright if I asked you another one of these later on?

hartnn (hartnn):

yes, sure :)

OpenStudy (anonymous):

Great! much obliged. :)